# Gibbs-preserving operations requiring infinite amount of quantum coherence

### v2

@@ A.4 Coherence cost for Gibbs-preserving operations
This motivates us to study the channel implementation cost introduced in Sec. A.1 in our setting, which corresponds to $\mathbb{F}=\mathbb{F}_{\rm inc}$, $\mathbb{O}_{1}=\mathbb{O}_{\rm GP}$, and $\mathbb{O}_{2}=\mathbb{O}_{\rm TO}$. For a coherence quantifier $R_{\mathbb{F}_{\rm inc}}$, we employ quantum Fisher information, the standard measure of coherence (and asymmetry in general), defined for an arbitrary state $\rho$ by
${\cal F}(\rho)=2\sum_{i,j}\frac{(p_{i}-p_{j})^{2}}{p_{i}+p_{j}}|\langle i|H|j\rangle|^{2}$ (23)
@@
$D_{F}(\rho,\sigma)=\sqrt{1-F(\rho,\sigma)^{2}},\quad F(\rho,\sigma)=\mathrm{Tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}.$ (26)
@@
A.5 Trade-off relation between symmetry, irreversibility, and quantum coherence
The main goal of this work is to evaluate the coherence cost in (27) for Gibbs-preserving Operations. In particular, we are interested in the fundamental limitations on the implementation of Gibbs-preserving Operations, which could be analyzed by obtaining lower bounds for ${\cal F}_{c}^{\epsilon}$. Lower bounds can never obtained by studying specific implementation protocols, and we thus need an approach that can put general restrictions on all feasible implementation strategies.
@@ A.5 Trade-off relation between symmetry, irreversibility, and quantum coherence
$\delta(\Lambda,\mathbb{P}):=\min_{{\cal R}:S^{\prime}\to S}\sqrt{\sum^{2}_{j=1}\frac{1}{2}D_{F}(\rho_{j},{\cal R}\circ\Lambda(\rho_{j}))^{2}}.$ (28)
@@
${\cal C}(\Lambda,\mathbb{P}):=\|\sqrt{\rho_{1}}(H_{S}-\Lambda^{\dagger}(H_{S^{\prime}}))\sqrt{\rho_{2}}\|_{2},$ (29)
where $\|O\|_{2}=\sqrt{\mathrm{Tr}(O^{\dagger}O)}$ is the Hilbert-Schmidt norm. We also define coherence cost for channel implementation by covariant operations
@@
$\frac{{\cal C}(\Lambda,\mathbb{P})}{\sqrt{{\cal F}_{c,{\rm cov}}(\Lambda)}+\Delta(H_{S})+\Delta(H_{S^{\prime}})}\leq\delta(\Lambda,\mathbb{P}),$ (31)
@@
We remark that the above relation can be extended from $\mathbb{P}$ to a general state ensemble $\Omega:=\{p_{j},\rho_{j}\}_{j}$ that may not be orthogonal to each other Tajima et al. (2022). The relation (31) is given as a unification between the Wigner-Araki-Yanase (WAY) theorems on quantum measurements Wigner (1952); Araki and Yanase (1960); Ozawa (2002a); Tajima and Nagaoka (2019); Kuramochi and Tajima (2023) and unitary gates Ozawa (2002b); Tajima et al. (2018, 2020); Tajima and Saito (2021) and the Eastin-Knill theorems on quantum error correcting codes Eastin and Knill (2009); Faist et al. (2020); Kubica and Demkowicz-Dobrzański (2021); Zhou et al. (2021); Yang et al. (2022); Tajima and Saito (2021). It also allows restrictions on the classical information recovery in the Hayden-Preskill thought experiments Hayden and Preskill (2007) imposed by the energy conservation Tajima et al. (2022) and extends the WAY theorem to various errors and disturbances of quantum measurements and the out-of-time-ordered correlators Emori and Tajima (2023). Here, we utilize this relation to obtain a lower bound for coherence cost for approximately implementing Gibbs-preserving Operations by Thermal Operations.
Appendix B Lower bound for pairwise reversible Gibbs-preserving Operations
(Proof of Theorem 2)
@@ Appendix B Lower bound for pairwise reversible Gibbs-preserving Operations
Let $\Lambda_{\epsilon}$ be a channel that approximates $\Lambda$ with error $\epsilon$, i.e., $D_{F}(\Lambda_{\epsilon},\Lambda)\leq\epsilon$ (recall (25)). We aim to apply Theorem S.1 to $\Lambda_{\epsilon}$ while expressing each term by the quantities in (32) relevant to the desired channel $\Lambda$ and the accuracy of implementation.
@@
$D_{F}({\cal R}\circ\Lambda_{\epsilon}(\rho),\rho)$ $\leq D_{F}({\cal R}\circ\Lambda_{\epsilon}(\rho),{\cal R}\circ\Lambda(\rho))+D_{F}({\cal R}\circ\Lambda(\rho),\rho)$ (33)
@@
$\delta(\Lambda_{\epsilon},\mathbb{P})\leq\sqrt{\frac{1}{2}D_{F}({\cal R}\circ\Lambda_{\epsilon}(\rho_{1}),\rho_{1})^{2}+\frac{1}{2}D_{F}({\cal R}\circ\Lambda_{\epsilon}(\rho_{2}),\rho_{2})^{2}}\leq\epsilon.$ (34)
@@
${\cal C}(\Lambda_{\epsilon},\mathbb{P})$ $=\|\sqrt{\rho_{1}}(H_{S}-\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}}))\sqrt{\rho_{2}}\|_{2}$ (35)
@@
$\|\sqrt{\rho_{1}}(\Lambda^{\dagger}(H_{S^{\prime}})-\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}}))\sqrt{\rho_{2}}\|_{2}$ $=\left\|\sum_{k,k^{\prime}}\sqrt{q_{k}}\sqrt{q_{k}^{\prime}}\psi_{k}(\Lambda^{\dagger}(H_{S^{\prime}})-\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}}))\phi_{k^{\prime}}\right\|_{2}$ (36)
@@
$|\langle\psi|\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\phi\rangle|$ $\leq\frac{1}{2}|\langle\psi|\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\phi\rangle+\langle\phi|\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\psi\rangle|$ (37)
@@
Together with (36), this particularly implies
$\|\sqrt{\rho_{1}}(\Lambda^{\dagger}(H_{S^{\prime}})-\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}}))\sqrt{\rho_{2}}\|_{2}$ $\leq\sqrt{\sum_{k,k^{\prime}}q_{k}q_{k^{\prime}}}2\Delta(H_{S^{\prime}})\epsilon$ (38)
@@
Combining this with (35), we get
${\cal C}(\Lambda_{\epsilon},\mathbb{P})\geq{\cal C}(\Lambda,\mathbb{P})-2\Delta(H_{S^{\prime}})\epsilon.$ (39)
We finally note that since $\mathbb{O}_{\rm TO}\subseteq\mathbb{O}_{\rm cov}$ as in (22), ${\cal F}_{c}^{\epsilon}(\Lambda)\geq{\cal F}_{c,\rm{cov}}^{\epsilon}(\Lambda)$ always holds. We conclude the proof by combining all these observations. Let us particularly take $\Lambda_{\epsilon}$ to be the optimal channel achieving the coherence cost with error $\epsilon$, i.e., $\Lambda_{\epsilon}\sim_{\epsilon}\Lambda$ and ${\cal F}_{c,{\rm cov}}(\Lambda_{\epsilon})={\cal F}_{c,{\rm cov}}^{\epsilon}(\Lambda)$. Then,
$\sqrt{{\cal F}_{c}^{\epsilon}(\Lambda)}$ $\geq\sqrt{{\cal F}_{c,{\rm cov}}^{\epsilon}(\Lambda)}$ (40)

### v3

@@ I.4 Coherence cost for Gibbs-preserving Operations
This motivates us to study the channel implementation cost introduced in Sec. I.1 in our setting, which corresponds to $\mathbb{F}=\mathbb{F}_{\rm inc}$, $\mathbb{O}_{1}=\mathbb{O}_{\rm GP}$, and $\mathbb{O}_{2}=\mathbb{O}_{\rm TO}$. For a coherence quantifier $R_{\mathbb{F}_{\rm inc}}$, we employ quantum Fisher information, the standard measure of coherence (and asymmetry in general), defined for an arbitrary state $\rho$ by
${\cal F}(\rho)=2\sum_{i,j}\frac{(p_{i}-p_{j})^{2}}{p_{i}+p_{j}}|\langle i|H|j\rangle|^{2}$ (S.23)
@@
$D_{F}(\Lambda_{1},\Lambda_{2})\coloneqq\max_{\rho}D_{F}(\operatorname{id}\otimes\Lambda_{1}(\rho),\operatorname{id}\otimes\Lambda_{2}(\rho))$ (S.25)
@@
$D_{F}(\rho,\sigma)=\sqrt{1-F(\rho,\sigma)^{2}},\quad F(\rho,\sigma)=\mathrm{Tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}.$ (S.26)
@@
I.5 Trade-off relation between symmetry, irreversibility, and quantum coherence
The main goal of this work is to evaluate the coherence cost in (S.27) for Gibbs-preserving Operations. In particular, we are interested in the fundamental limitations on the implementation of Gibbs-preserving Operations, which could be analyzed by obtaining lower bounds for ${\cal F}_{c}^{\epsilon}$. Lower bounds can never obtained by studying specific implementation protocols, and we thus need an approach that can put general restrictions on all feasible implementation strategies.
@@ I.5 Trade-off relation between symmetry, irreversibility, and quantum coherence
$\delta(\Lambda,\mathbb{P}):=\min_{{\cal R}:S^{\prime}\to S}\sqrt{\sum^{2}_{j=1}\frac{1}{2}D_{F}(\rho_{j},{\cal R}\circ\Lambda(\rho_{j}))^{2}}.$ (S.28)
@@
${\cal C}(\Lambda,\mathbb{P}):=\|\sqrt{\rho_{1}}(H_{S}-\Lambda^{\dagger}(H_{S^{\prime}}))\sqrt{\rho_{2}}\|_{2},$ (S.29)
where $\|O\|_{2}=\sqrt{\mathrm{Tr}(O^{\dagger}O)}$ is the Hilbert-Schmidt norm. We here remark that ${\cal C}(\Lambda,\mathbb{P})$ is well-defined even when $\mathbb{P}$ is not a perfectly reversible pair for $\Lambda$. The quantity ${\cal C}(\Lambda,\mathbb{P})$ is well-defined for an arbitrary CPTP map $\Lambda$ and an arbitrary orthogonal state pair =$\mathbb{P}=\{\rho_{1},\rho_{2}\}$.
We also define coherence cost for channel implementation by covariant operations
@@
$\frac{{\cal C}(\Lambda,\mathbb{P})}{\sqrt{{\cal F}_{c,{\rm cov}}(\Lambda)}+\Delta(H_{S})+\Delta(H_{S^{\prime}})}\leq\delta(\Lambda,\mathbb{P}),$ (S.31)
@@
We remark that the above relation can be extended from $\mathbb{P}$ to a general state ensemble $\Omega:=\{p_{j},\rho_{j}\}_{j}$—where $\{p_{j}\}_{j}$ is probability distribution and $\{\rho_{j}\}_{j}$ are states—that may not be orthogonal to each other Tajima et al. (2022). The relation (S.31) is given as a unification between the Wigner-Araki-Yanase (WAY) theorems on quantum measurements Wigner (1952); Araki and Yanase (1960); Ozawa (2002a); Tajima and Nagaoka (2019); Kuramochi and Tajima (2023) and unitary gates Ozawa (2002b); Tajima et al. (2018, 2020); Tajima and Saito (2021) and the Eastin-Knill theorems on quantum error correcting codes Eastin and Knill (2009); Faist et al. (2020); Kubica and Demkowicz-Dobrzański (2021); Zhou et al. (2021); Yang et al. (2022); Tajima and Saito (2021). It also allows restrictions on the classical information recovery in the Hayden-Preskill thought experiments Hayden and Preskill (2007) imposed by the energy conservation Tajima et al. (2022) and extends the WAY theorem to various errors and disturbances of quantum measurements and the out-of-time-ordered correlators Emori and Tajima (2023). Here, we utilize this relation to obtain a lower bound for coherence cost for approximately implementing Gibbs-preserving Operations by Thermal Operations.
Next, we discuss the physical meaning of ${\cal C}$. To provide an intuitive understanding, the following inequality is useful Tajima et al. (2022):
$\left(\min_{k}\lambda^{\min}_{>0}(\rho_{k})\right)\frac{{\cal C}_{F}(\Lambda,\mathbb{P})}{2}\leq{\cal C}(\Lambda,\mathbb{P})^{2}\leq\frac{{\cal C}_{F}(\Lambda,\mathbb{P})}{2}.$ (S.32)
Here, $\{\rho_{k}\}_{k=1,2}$ are the states in $\mathbb{P}$, and $\lambda^{\min}_{>0}(\rho_{k})$ is the minimum non-zero eigenvalue of $\rho_{k}$. The quantity ${\cal C}_{F}(\Lambda,\mathbb{P})$ represents the degree of convexity of the Fisher information with respect to the work operator $Y\coloneqq H_{S}-\Lambda^{\dagger}(H_{S^{\prime}})$ as
${\cal C}_{F}(\Lambda,\mathbb{P})$ $:=\sum_{k=1,2}p_{k}{\cal F}(\rho_{k};Y)-{\cal F}\left(\sum_{k=1,2}p_{k}\rho_{k};Y\right),$ (S.33)
$p_{k}$ $:=\frac{1}{2}\enskip(k=1,2).$ (S.34)
Here we made the dependence on the observable $Y$ explicit. Therefore, ${\cal C}$ essentially characterizes the extent to which the coherence (Fisher information) of $Y$ diminishes when the states $\rho_{1}$ and $\rho_{2}$ are probabilistically mixed. Importantly, when $\mathbb{P}$ consists of a pair of orthogonal pure states, as in the scenarios discussed in this letter, the inequality (S.32) reduces to ${\cal C}^{2}={\cal C}_{F}/2$—meaning that the quantity ${\cal C}$ is characterized by the convexity of Fisher information. We also remark that (S.32) can be extended to the case involving a general test ensemble $\Omega=\{p_{j},\rho_{j}\}_{j}$ as long as $\{\rho_{j}\}_{j}$ are orthogonal to each other.
The physical meaning of ${\cal C}_{F}$ (and thus ${\cal C}$) becomes clear when we note that ${\cal C}_{F}$ is analogous to the Holevo quantity $\chi:=S(\sum_{j}p_{j}\rho_{j})-\sum_{j}p_{j}S(\rho_{j})$, with entropy being replaced by Fisher information with respect to $Y$. Holevo’s $\chi$ quantifies the information loss in terms of entropy when states are probabilistically mixed. Consequently, ${\cal C}_{F}$ can be understood as an indicator of the resource loss in terms of Fisher information with respect to $Y$ when the states in $\mathbb{P}$ are probabilistically mixed.
II Lower bound for pairwise reversible Gibbs-preserving Operations (Proof of Theorem 1)
II.1 Proof of Theorem 1
@@
$\sqrt{{\cal F}^{\epsilon}_{c}(\Lambda)}\geq\frac{{\cal C}(\Lambda,\mathbb{P})}{\epsilon}-\Delta(H_{S})-3\Delta(H_{S^{\prime}}),$ (S.35)
@@
Let $\Lambda_{\epsilon}$ be a channel that approximates $\Lambda$ with error $\epsilon$, i.e., $D_{F}(\Lambda_{\epsilon},\Lambda)\leq\epsilon$ (recall (S.25)). We aim to apply Theorem S.1 to $\Lambda_{\epsilon}$ while expressing each term by the quantities in (S.35) relevant to the desired channel $\Lambda$ and the accuracy of implementation.
@@
$D_{F}({\cal R}\circ\Lambda_{\epsilon}(\rho),\rho)$ $\leq D_{F}({\cal R}\circ\Lambda_{\epsilon}(\rho),{\cal R}\circ\Lambda(\rho))+D_{F}({\cal R}\circ\Lambda(\rho),\rho)$ (S.36)
@@
$\delta(\Lambda_{\epsilon},\mathbb{P})\leq\sqrt{\frac{1}{2}D_{F}({\cal R}\circ\Lambda_{\epsilon}(\rho_{1}),\rho_{1})^{2}+\frac{1}{2}D_{F}({\cal R}\circ\Lambda_{\epsilon}(\rho_{2}),\rho_{2})^{2}}\leq\epsilon.$ (S.37)
@@
${\cal C}(\Lambda_{\epsilon},\mathbb{P})$ $=\|\sqrt{\rho_{1}}(H_{S}-\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}}))\sqrt{\rho_{2}}\|_{2}$ (S.38)
@@
$\|\sqrt{\rho_{1}}(\Lambda^{\dagger}(H_{S^{\prime}})-\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}}))\sqrt{\rho_{2}}\|_{2}$ $=\left\|\sum_{k,k^{\prime}}\sqrt{q_{k}}\sqrt{q_{k}^{\prime}}\psi_{k}(\Lambda^{\dagger}(H_{S^{\prime}})-\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}}))\phi_{k^{\prime}}\right\|_{2}$ (S.39)
@@
$|\langle\psi|\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\phi\rangle|$ $\leq\frac{1}{2}|\langle\psi|\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\phi\rangle+\langle\phi|\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}})-\Lambda^{\dagger}(H_{S^{\prime}})|\psi\rangle|$ (S.40)
@@
Together with (S.39), this particularly implies
$\|\sqrt{\rho_{1}}(\Lambda^{\dagger}(H_{S^{\prime}})-\Lambda_{\epsilon}^{\dagger}(H_{S^{\prime}}))\sqrt{\rho_{2}}\|_{2}$ $\leq\sqrt{\sum_{k,k^{\prime}}q_{k}q_{k^{\prime}}}2\Delta(H_{S^{\prime}})\epsilon$ (S.41)
@@
Combining this with (S.38), we get
${\cal C}(\Lambda_{\epsilon},\mathbb{P})\geq{\cal C}(\Lambda,\mathbb{P})-2\Delta(H_{S^{\prime}})\epsilon.$ (S.42)
We finally note that since $\mathbb{O}_{\rm TO}\subseteq\mathbb{O}_{\rm cov}$ as in (S.22), ${\cal F}_{c}^{\epsilon}(\Lambda)\geq{\cal F}_{c,\rm{cov}}^{\epsilon}(\Lambda)$ always holds. We conclude the proof by combining all these observations. Let us particularly take $\Lambda_{\epsilon}$ to be the optimal channel achieving the coherence cost with error $\epsilon$, i.e., $\Lambda_{\epsilon}\sim_{\epsilon}\Lambda$ and ${\cal F}_{c,{\rm cov}}(\Lambda_{\epsilon})={\cal F}_{c,{\rm cov}}^{\epsilon}(\Lambda)$. Then,
$\sqrt{{\cal F}_{c}^{\epsilon}(\Lambda)}$ $\geq\sqrt{{\cal F}_{c,{\rm cov}}^{\epsilon}(\Lambda)}$ (S.43)